OpenStudy (anonymous):

@satellite73 Did I get this problem right? What is the oxidation number of sulfur in the S2O8^2- ion? 2s+8(-2)=-2 2s-16=-2 2s=14 s=7

OpenStudy (anonymous):

@agent0smith Did I get this problem right?

OpenStudy (anonymous):

@mathstudent55 Did I get this problem right?

OpenStudy (photon336):

yes you are correct @badwolf1646 Whenever you have an ion the sum of the oxidation states equals the ion. so in this case we have -2, so the sum of oxidation states must equal -2. we have 8 atoms of oxygen so let's make an equation \[2S+8O = -2 \] we know the oxidation state of oxygen is -2 so let's re-arrange for sulfur \[\frac{ -2-(8O) }{ 2 } = S\] Now, we plug in our value for oxygen -2 \[\frac{ -2-8(-2) }{ 2 } = \frac{ -2+16 }{ 2 } = \frac{ 14 }{ 2 } = 7 \] so each sulfur atom has an oxidation state of 7